OpenStudy (anonymous):

How would I factor 3x^3 - x^2 + 3x - 1?

OpenStudy (anonymous):

I could be wrong, but you could separate it by doing this\[(3x ^{3}-x ^{2})+(3x-1) \]

OpenStudy (anonymous):

So then factor a 3x^2 from first parenthesis and a 3 from the second\[3x ^{2}(x-1)+3(x-(\frac{ 1 }{ 3 }))=0\]So just find x when x=0\[3x ^{2}=0\]and\[x-1=0\]lastly\[x-\frac{ 1 }{ 3 }=0\]

OpenStudy (anonymous):

If I'm wrong, sorry its been about 5 good years when I actually had a good teacher who taught me math.

OpenStudy (anonymous):

Thank you for all your help. I have been working on this one problem for about an hour.